OpenStudy (anonymous):

Which student factored the expression x^3 + 2^x - 3x^2 - 6 correctly based on their answer below? Jane: (x + 3)(x^2 + 2) Mariah: (x - 3)(x^2 + 2) Roger: (x - 3)(x^2 - 2) Nethan: (x + 3) x^2 - 2) Jane Mariah Roger Nethan

OpenStudy (solomonzelman):

It really seems impossible to factor it since it contains \(\normalsize\color{black}{ 2^x }\) .

OpenStudy (lochana):

i think it's 2x right?, not 2^x and answer is mariah

OpenStudy (solomonzelman):

Yes, if it is \(\normalsize\color{black}{ 2x }\), and not \(\normalsize\color{black}{ 2^x }\) you can factor it (one of the factors is `x-1` ) \(\normalsize\color{black}{ ✂≠ θ∠∆∓⇒ \rm { }✓ }\)

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

Which shows 632 - 372 being evaluated using the difference of squares method? 632 - 372 = (3969 + 1369)(3969 - 1369) = 13,878,800 632 - 372 = 3969 - 1369 = 2,600 632 - 372 = (63 - 37)2 = 262 = 676 632 - 372 = (63 + 37)(63 - 37) = (100)(26) = 2,600